The mean IQ is 100 with a standard deviation of 15. The Z-score of an IQ of 117 is 

1
Expert's answer
2021-05-26T17:37:45-0400

"X\sim (\mu, \sigma^2)"

Given "\mu=100, \sigma=15"


"Z=\dfrac{X-\mu}{\sigma}=\dfrac{117-100}{15}=\dfrac{17}{15}\approx1.133333"

"Z=\dfrac{17}{15}\approx1.133333"